OpenStudy (anonymous):

Verify that the equation is an identity. sin^2(theta/ 2 ) = tan (theta ) - sin ( theta ) / 2 tan ( theta)

OpenStudy (anonymous):

\[\sin ^{2}\frac{ \theta }{ 2} = \frac{ \tan \theta - \sin \theta}{ 2 \tan \theta}\]

OpenStudy (fibonaccichick666):

so, using known identities, you have to turn the left side into the right